please help me solve this equation: c/2=7
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Help you do what?

Answer by stanbon(75887)

c/2=7
---
Multiply both sides by 2 to get:
c = 2*7 = 14
==============
Cheers,
Stan H.
==============